FEL Upgrade Project Weekly Brief
January 29-February 2, 2001



FEL operations delivered beam for experiments with pulsed laser deposition, carbon nanotube production , metal ablation and beam characterization.


The FEL Upgrade Project monthly reports for December 2000 were distributed to the DOE and ONR program offices this week.  We are pleased to announce that the DOE Jefferson Lab Site Office received the funding for the second phase of the FEL Upgrade project covering the period from Feb. 1,2001 until Sept. 30,2002.

WBS 3 (Beam Physics):

Work began on beamline layout and beam optics design for the Revision 1.1 UV bypass.

WBS 4 (Injector):

Gun HVPS - Discussions are ongoing with Glassman concerning the need for an additional multiplier stack.  R.Walker's  circuit simulation did not indicate a significant change in the dissipation of the HV capacitors when the current load was
changed from 5 to 10 ma.  We are  requesting their analysis of the problem in writing.

WBS 6 (RF):

Zone 3 - No effort this week.

Zone 4 - The zone is operating in HV of 11 kV to process and condition a klystron.  At present the Crowbar fires
when the HVPS current reaches 5.8 amps.  We operate approximately at 11 amps and the HVPS should be capable of 15 amps, so we have a ways to go.  RF Control Modules were installed and workable Arc Detector heads are to be available next week for RF testing into shorts.

Injector RF - The 100 kW klystrons are on schedule for April delivery. Their HVPS upgrade is scheduled for the following year.  We will test them with the present HVPS at reduced power levels. PSC is reviewing our latest inquiry concerning replacing the 3 phase variac with a tapped transformer and a soft-start controller.

Operations - RF systems had and solved 5 major problems this week:

1.  HVPS Inj. Cavity 4 - Variac fuses opened in one phase and burned up a fuse board.  Fuses were replaced, but the
     burned board was not.
2.  HVPS Inj. Cavity 4 - Variac fuses for the same phase opened again after 6 hours of operation.  Fuses and the missing
     fuse board were replaced.
3.  HVPS Inj. Cavity 4 - Variac fuses opened for a different phase.  All Variac brushes were replaced and the coils
4.  The fuse for the -18 volt power supply opened for Zone 1.  Problem was solved by cutting out shorted capacitors on
     the back of the RF Control Module Crate.
5.  HVPS Inj. Cavity 4 - The Klystron Current trip level drift level drifted lower and the HVPS wouldn't stay on.  The
     HVPS was turned on manually to a slightly lower voltage and current for a shift and the trip level was reset the next
     day  for automatic operation.

WBS 8 (I&C):

This week the surface mount version of the optical beam position monitor was completed and sent out for manufacturing.
The locking software has not yet been tried out due to limited availability of the FEL this week.  Meetings were held to discuss the transition from the operating FEL to an "empty" vault and to the IR upgrade.  There was a visit to the machine shop working on the 3" BPMs and viewers all parts are on schedule and looking great!  The shop has 3 CNC machines cranking out the pieces. Other efforts included support of the user runs this week.

WBS 9 (Transport):

Injector Dipoles (DU/DV)
o DULY Research continues to model the small Injector dipole.  They are bench marking their RADIA magnetic model
   program vs. our TOSCA program.
Optical Chicane Dipole (DW)
o We reviewed the design of this prototype dipole with the engineer and technician in Magnet Measurement to catch any
    potential problems.  We found a better way to build the field clams and incorporated their suggestions in the details.
    The details are now ready for tolerance analysis and checking.
Arc Dipoles (GY, GX, GQ)
o AES incorporated the GG correctors extra steel and the better way of building the field clamps into their drawings of
   the 180° Dipole (DY).  They are now starting to layout the coil's leads, incorporating the 4-in-hand winding.
o AES started their magnetic modeling of the GY.
o We determined the shape and size of the First Arc Bend (GX) with the provisions for UV extraction to transmit to AES
   for their future layout.
QX (3.125" Quad)
o Measurement Probe: The coil support plate bow was minimized by cutting 1/16 inch slots and was then glued to the
   half tube to form the final section.  Deflection was measured and found to be an acceptable 0.0004 inch.  We sent it
   back to the machining vendor for final machining of the coil slots.
o Prices for steel were received from Gary Steel and are being evaluated
o We had a brainstorming session to integrate the quadrupole yoke high tolerance inner cut out with the Magnet Test
   Stand, the fiducials for alignment and mounting to girders.
Sextupole with H/V Correctors
o Robin Wines is optimizing the magnetic model of the Sextupole before trying to introduce the superposed octupole

WBS 10 (Wiggler):

Magnetic Measurements:  C. Hutton continued on a rail design for Hall Probe measurements.
o Assembly:  The mounts for the Dispersion Section were placed on the stand for the Wiggler.
o Dispersion Section: Coils were shipped to the primary assembly vendor.
o Steve Benson will do the magnetic analysis using RADIA to find what correction windings and end pole modifications
   are necessary to add a field integral of half of a standard pole to the end poles.  Those regions are now running at 3% of
   a pole and are causing the bias in the fields from pole to pole reported earlier.

WBS 11 (Optics):

The prints we received as part of the AES subcontract were imported into the I-DEAS CAD system, and made some of the
modifications we identified during our review.  We are reviewing the check prints.  The majority of our time was spent supporting operations, details follow.

The OBPM system works well; and we've added optical hardware to the OBPM in the optics control room (OCR), the first
installed, to bring it to the same level of performance as the later version.  It will be aligned Monday, when we use the optical
transport system.


Operations delivered beam for the nanotubes group, for PLD experiments, for metal ablation studies, and for FEL optical mode and beam quality characterization studies this week.  We lost some time due to drive laser problems earlier in the
week, rf problems mid week and a temporary low helium level in the injector quarter unit yesterday.  Mr. Murphy seems to have occupied the FEL building this week.

When the drive laser was started Monday morning, we immediately discovered there was no IR output.  Investigation showed that a cooling water leak was spoiling the gain.  Surmising that a Teflon seal on the laser rod had failed, the rod assembly was removed and the spare installed. It also leaked. Some time was spent before it was determined that it was a different seal that had failed. That was replaced and the drive laser was working again at 17:30 Tuesday.


Privacy and Security Notice